OpenStudy (anonymous):

Identify the slope and y-intercept for each of the linear equations. a. 2x + 4y = 14 b. -12x + 3y = -16 c. x - 15y = 45

OpenStudy (bobbyleary):

The best way to find your answers is to put them in the form y = mx + b, where m is your slope and b is your y-intercept. For a: 4y = - 2x + 14 y = (-1/2)x + (14/4) This means the slope is (-1/2) and the y-intercept is (14/4) = (7/2) Try the rest and post your solutions.

OpenStudy (anonymous):

a) y=-1/2 x+7/2 so, slope m=-1/2 y-intercept (x=0) y=7/2 try the rest on your own, please